Implement the following multiple output combinational logic circuit using a 4 line to 16-line decoder: F1= ∑m (0, 1, 4, 7, 12, 14, 15) F3= ∑m (2, 3, 7, 8, 10) F2= ∑m (1, 3, 6, 9, 12) F4= ∑m (1, 3, 5)


1
Expert's answer
2021-05-30T15:49:37-0400

Implementation:

"F1=\sum m (0,1,4,7,12,14,15)"





"F2=\sum m(1,3,6,9,12)"




"F3=\sum m (2,3,7,8,10)"





"F4=\sum m(1,3,5)"